# Eclipsing Stellar Flare on the Demon Star Algol Binary System Observed during the MAXI-NICER Follow-up Campaign in 2018

### v1

@@ Abstract
facilities: MAXI (Matsuoka et al., 2009), NICER (Gendreau et al., 2016), TESS (Ricker et al., 2014)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), HEASoft, Xspec (Arnaud, 1996), SKIRT (Baes et al., 2003)
@@ 1 Introduction
2 Observations and Data Reduction
@@ 3 Analysis and Results
4 Discussion
Here, we estimate the location and height of the MAXI-NICER flare to explain the observations described in Section 3. Since the Algol binary system has a short orbital period of $P_{\mathrm{{orb}}}\simeq 2.867$ d, and the two stars are thought to be tidally locked, the stellar rotation period of Algol B is assumed to be the same as $P_{\mathrm{{orb}}}$. Then, we approximated the flare loop as a cylinder structure attached and perpendicular to the stellar surface, the interior of which is filled with a uniform CIE plasma. Figure 5 shows the two examples identified in our numerical study. These examples are suggested based on the following observational requirements.
@@ 4 Discussion
from Shibata & Yokoyama (2002), where $L$, EMEM, $n_{0}$, and $T$ are the flare loop length, the peak emission measure, the preflare electron density, and the peak temperature, respectively. This equation is commonly used when calculating $L$ from the X-ray spectroscopy of stellar flares (Raassen, Mitra-Kraev, & Güdel, 2007; Namekata et al., 2017; Notsu et al., 2024). The flare peak parameters at the MAXI scan are $EM=3.7\times 10^{55}~\mathrm{cm^{-3}}$ and $T=6.0~\mathrm{keV}=7.0\times 10^{7}~\mathrm{K}$ ($T$ is fixed to a constant value when calculating the $EM$ in Section 3). Since we do not know $n_{0}$ from the observations, we assumed a typical range of $n_{0}=10^{10-12}~\mathrm{cm^{-3}}$ from the X-ray grating observations in Ness et al. (2004). As a result, we estimated $L=(1.0-6.2)\times 10^{11}~\mathrm{cm}=0.4-2.6R_{B}$. If $H$ is thought to be half the flare loop length, then $H=(0.5-3.1)\times 10^{11}~\mathrm{cm}=0.2-1.3R_{B}$ is derived. The loop height from the eclipse geometry (in the above) is in the range estimated here, and this suggests the both estimates can be consistent.
@@ 5 Summary
We sincerely thank Dr. Jerzy M Kreiner of Mt. Suhora Astronomical Observatory for kindly providing the O-C data (Kreiner, 2004) of Algol used in this study. Some/all of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/5dyy-t366 (catalog doi: 10.17909/5dyy-t366). The NICER analysis software and data calibration were provided by the NASA NICER mission and the Astrophysics Explorers Program. This research has made use of the MAXI data provided by RIKEN, JAXA and the MAXI team. W.I was supported by JSPS KAKENHI JP24K00673. T.E. was supported by the JST, Japan grant number JPMJFR202O (Sohatsu). Y.N. acknowledges the support from the NASA ADAP award program Number 80NSSC21K0632, the NASA TESS Cycle 6 Program 80NSSC24K0493, and the NASA NICER Cycle 6 Program 80NSSC24K1194. The work of K.H. is supported by NASA under award number 80GSFC24M0006. To verify whether the observed spectrum during the eclipse can be reproduced by these configurations, we tried 3D radiation transfer simulations with the Stellar Kinematics Including Radiative Transfer (SKIRT) toolkit (Baes et al., 2003). It is a Monte Carlo simulation to emulate scattering, absorption, and emission by the surrounding medium around the X-ray source. Using this, we tried to simulate the observed eclipsing spectrum by assuming the incident non-eclipsing spectral energy distribution (SED). The stellar positions of the Algol binary system are assumed to be at mid-eclipse (Figure 5 panels b1, b2). We approximate the emission region by placing three point sources at the base, middle, and top of the flare loop in the cylinder shape, with each emitting intensity at one-third of the total non-eclipse incident X-ray flux. As the incident SED setup, we assumed the NICER best-fit spectral model observed before and after the eclipse (MJD 58304.34 and 58304.67, ObsID 1200260102). Under these assumptions, we confirmed that the normalization of the eclipsing spectra is reduced by 70% from the incident model in both configurations. The input model assuming the post-eclipse showed a better match to the observed data than the model of the pre-eclipse case. The simulated spectrum assuming the post-eclipse model as the input stays within 0.8–1.2 level of the observed spectral shape in the 0.3–3.0 keV. Thus, we confirmed the assumed configuration approximately reproduce the observed spectrum.

### v2

@@ Abstract
facilities: MAXI (Matsuoka et al., 2009), NICER (Gendreau et al., 2016), TESS (Ricker et al., 2014)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), HEASoft, Xspec (Arnaud, 1996), SKIRT (Baes et al., 2003; Vander et al., 2023)
@@ I Introduction
II Observations and Data Reduction
@@ III Analysis and Results
IV Discussion
Here, we estimate the location and height of the MAXI-NICER flare to explain the observations described in Section III. Since the Algol binary system has a short orbital period of $P_{\mathrm{{orb}}}\simeq 2.867$ d, and the two stars are thought to be tidally locked, the stellar rotation period of Algol B is assumed to be the same as $P_{\mathrm{{orb}}}$. Then, we approximated the flare loop as a cylinder structure attached and perpendicular to the stellar surface, the interior of which is filled with a uniform CIE plasma. Figure 5 shows the two examples identified in our numerical study. These examples are suggested based on the following observational requirements.
@@ IV Discussion
from Shibata & Yokoyama (2002), where $L$, EMEM, $n_{0}$, and $T$ are the flare loop length, the peak emission measure, the preflare electron density, and the peak temperature, respectively. This equation is commonly used when calculating $L$ from the X-ray spectroscopy of stellar flares (Raassen, Mitra-Kraev, & Güdel, 2007; Namekata et al., 2017; Notsu et al., 2024). The flare peak parameters at the MAXI scan are $EM=3.7\times 10^{55}~\mathrm{cm^{-3}}$ and $T=6.0~\mathrm{keV}=7.0\times 10^{7}~\mathrm{K}$ ($T$ is fixed to a constant value when calculating the $EM$ in Section III). Since we do not know $n_{0}$ from the observations, we assumed a typical range of $n_{0}=10^{10-12}~\mathrm{cm^{-3}}$ from the X-ray grating observations in Ness et al. (2004). As a result, we estimated $L=(1.0-6.2)\times 10^{11}~\mathrm{cm}=0.4-2.6R_{B}$. If $H$ is thought to be half the flare loop length, then $H=(0.5-3.1)\times 10^{11}~\mathrm{cm}=0.2-1.3R_{B}$ is derived. The loop height from the eclipse geometry (in the above) is in the range estimated here, and this suggests the both estimates can be consistent.
@@ V Summary
We sincerely thank Dr. Jerzy M Kreiner of Mt. Suhora Astronomical Observatory for kindly providing the O-C data (Kreiner, 2004) of Algol used in this study. Some/all of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/5dyy-t366 (catalog doi: 10.17909/5dyy-t366). The NICER analysis software and data calibration were provided by the NASA NICER mission and the Astrophysics Explorers Program. This research has made use of the MAXI data provided by RIKEN, JAXA and the MAXI team. W.I was supported by JSPS KAKENHI JP24K00673. T.E. was supported by the JST, Japan grant number JPMJFR202O (Sohatsu). Y.N. acknowledges the support from the NASA ADAP award program Number 80NSSC21K0632, the NASA TESS Cycle 6 Program 80NSSC24K0493, and the NASA NICER Cycle 6 Program 80NSSC24K1194. The work of K.H. is supported by NASA under award number 80GSFC24M0006.
To verify whether the observed spectrum during the eclipse can be reproduced by these configurations, we tried 3D radiation transfer simulations with the Stellar Kinematics Including Radiative Transfer (SKIRT) toolkit (Baes et al., 2003; Vander et al., 2023). It is a Monte Carlo simulation to emulate scattering, absorption, and emission by the surrounding medium around the X-ray source. Using this, we tried to simulate the observed eclipsing spectrum by assuming the incident non-eclipsing spectral energy distribution (SED). The stellar positions of the Algol binary system are assumed to be at mid-eclipse (Figure 5 panels b1, b2). We approximate the emission region by placing three point sources at the base, middle, and top of the flare loop in the cylinder shape, with each emitting intensity at one-third of the total non-eclipse incident X-ray flux. As the incident SED setup, we assumed the NICER best-fit spectral model observed before and after the eclipse (MJD 58304.34 and 58304.67, ObsID 1200260102). Under these assumptions, we confirmed that the normalization of the eclipsing spectra is reduced by 70% from the incident model in both configurations. The input model assuming the post-eclipse showed a better match to the observed data than the model of the pre-eclipse case. The simulated spectrum assuming the post-eclipse model as the input stays within 0.8–1.2 level of the observed spectral shape in the 0.3–3.0 keV. Thus, we confirmed the assumed configuration approximately reproduce the observed spectrum.